A car is initially travelling at 5.0ms-1. It then speeds up, accelerating at 1.2 ms-2 for 4.0 seconds. What is the cars final velocity?
1
Expert's answer
2019-04-11T09:34:13-0400

We can find the car's final velocity from the kinematic equation:


"v_f = v_i + at,"

here, "v_i = 5.0 ms^{-1}" is the initial velocity of the car, "v_f" is the final velocity of the car, "a = 1.2 ms^{-2}" is the acceleration of the car and "t = 4.0 s" is the time during which the car accelerates.

Then, we get:


"v_f = v_i + at = 5.0 \dfrac{m}{s} + 1.2 \dfrac{m}{s^2} \cdot 4.0 s = 9.8 \dfrac{m}{s}."

Answer:

"v_f = 9.8 \dfrac{m}{s}."